# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ 4.3 MdFOCuS(0): a Simple Online Algorithm
In this section we propose an algorithm to compute the GLR statistics at every time step $n$ for a known and unknown $\eta_{1}.$ In both cases, based on Theorem 5 we only need to consider the changepoints $\tau$ such that $P(\tau)$ is on the hull of $\left\{P(\tau)\right\}_{\tau\in\{1,\dots,n-1\}}.$
@@
Let us call $\mathcal{T}_{n}$ the set of index $\tau$ such that $P(\tau)=(\tau,\sum_{t=1}^{\tau}x_{t})$ is on the convex hull of all $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$. Restating Theorem 4, the changepoint $\hat{\tau}$ maximising the likelihood (6) (i.e $\hat{\tau}(\eta_{1})$ if $\eta_{1}$ is known or $\hat{\tau}(.)$ if it is not) is in $\mathcal{T}_{n}$. Hence, computationally speaking, at step $n$ we only need to compute the likelihood (and likelihood ratio statistics) of changepoints $\tau$ in $\mathcal{T}_{n}$. Furthermore, as already explained after Theorem 5 we can prune all $\tau$ that are not in $\mathcal{T}_{n}$ as they will never be in $\mathcal{T}_{n^{\prime}}$ for any $n^{\prime}\geq n$.
@@
Computing the convex hull from scratch at every time step $n$ is inefficient: it would require at least $\mathcal{O}(n)$ operations and lead to an overall complexity at least quadratic in the number of data points. Ideally, we would like to update the hull using an efficient online or dynamic convex hull algorithm as the one in [60]. There exists implementations of such an algorithm in dimension $p=2$ [61] and $p=3$ [62] However, out of simplicity, here we rely on an iterative use of the offline QuickHull algorithm [48] that also works for larger dimensions and is available in the geometry R package [59].
@@
In detail, for any $n^{\prime}>n$ we have $\mathcal{T}_{n^{\prime}}\subset\mathcal{T}_{n}\cup\{n,\ldots,n^{\prime}-1\}$. Therefore if we have already computed $\mathcal{T}_{n}$, with QuickHull, when we compute $\mathcal{T}_{n^{\prime}}$ we need not restart from scratch but simply apply QuickHull to the points $P(\tau)$ with $\tau$ in $\mathcal{T}_{n}\cup\{n,\ldots,n^{\prime}-1\}$. A first idea would be to apply the QuickHull algorithm at every time step $n,n+1,\ldots$, that is compute $\mathcal{T}_{n}$ as the hull of all $P(\tau)$ with $\tau$ in $\mathcal{T}_{n-1}\cup\{n\}$. In practice, we discovered that running the QuickHull algorithm intermittently, rather than at every iteration, resulted in faster run times. Specifically, we maintain a set of linearly increasing indices, denoted as $\mathcal{T}^{\prime}_{n}$, and only periodically execute QuickHull to reconstruct the true $\mathcal{T}_{n}$. More precisely, to decide when to run QuickHull we update a maximum size variable for $\mathcal{T}^{\prime}_{n}$ (denoted $maxSize$). Then, at step $n+1$:
  • either the size of $\mathcal{T}^{\prime}_{n}$, $|\mathcal{T}^{\prime}_{n}|$, is smaller than $maxSize$ in that case we set $\mathcal{T}^{\prime}_{n+1}$ to $\mathcal{T}^{\prime}_{n}\cup\{n\}$,
  • or it is larger and in that case we compute $\mathcal{T}^{\prime}_{n+1}$ as the convex hull of all $P(\tau)$ with $\tau$ in $\mathcal{T}^{\prime}_{n}\cup\{n\}$.
@@
In the later case we also update $maxSize$ to $maxSize=\lfloor\alpha|\mathcal{T}^{\prime}_{n+1}|+\beta\rfloor$, with $\alpha\geq 1$ and $\beta\geq 0$. It is important to note that, in this way, we always retain a few more possible changepoints than necessary: $\mathcal{T}_{n+1}\subseteq\mathcal{T}^{\prime}_{n+1}$.
@@
Taking $\beta=1$ and $\alpha=1$, we run the QuickHull algorithm at every time step. Increasing the value of $\alpha$ allows to run it less often at the cost of having a longer list of indices $\mathcal{T}^{\prime}_{n}$. In Figure 11 we considered different value for $\alpha$ with $\beta=1$ and found that a value of $\alpha=2$ gives good practical performances (see Supplementary Material E.1).
@@
A formal description of our procedure, called MdFOCuS if $\eta_{1}$ is unknown and MdFOCuS0 if it is, is available in Algorithm 1. The only difference between MdFOCuS and MdFOCuS0 variants is in the calculation of the likelihood (lines 9 and 10). Therefore the two variants store the same number of candidate changepoints and have the same time and memory complexities. In detail, computing the full likelihood ratio statistics for any particular changepoint can be done in $\mathcal{O}(p)$ and therefore lines 9 and 10 in Algorithm 1 can be done $\mathcal{O}(|\mathcal{T}|p).$ Recovering the maximum is then in $\mathcal{O}(|\mathcal{T}|).$ Overall, this is $\mathcal{O}(|\mathcal{T}|p).$
Algorithm 1 MdFOCuS(0) algorithm
1:Input 1: $p$-variate independent time series $\{x_{t}\}_{t=1,2,\dots}$
2:Input 2: threshold $thrs$ and $\eta_{1}$ if known pre-change parameter
3:Input 3: pruning index $maxSize>p+1$, the parameters $\alpha\geq 1$, $\beta\geq 0$.
4:Output: stopping time $n$ and maximum likelihood change : $\hat{\tau}(.)$ or $\hat{\tau}(\eta_{1})$
5:$logLikelihoodRatio\leftarrow-\infty,\quad\mathcal{T}\leftarrow\emptyset,\quad n\leftarrow 0$
6:while ($logLikelihoodRatio<thrs$do
7:     $n\leftarrow n+1$
8:     $\mathcal{T}\leftarrow\mathcal{T}\cup\{n-1\}$
9:     $maxLogLikelihood\leftarrow\max_{\tau\in\mathcal{T}}\{\max_{\eta_{1},\eta_{2}}\ell_{\tau,n}(\eta_{1},\eta_{2})\}$triangleright $\eta_{1}$ is fixed in MdFOCuS0
10:     $logLikelihoodRatio\leftarrow maxLogLikelihood-\max_{\eta_{1},\eta_{2}}\ell_{n,n}(\eta_{1},\eta_{2})$triangleright $\eta_{1}$ is fixed in MdFOCuS0
11:     if ($|\mathcal{T}|>maxSize$then
@@
12:         $\mathcal{T}\leftarrow\textsc{Quickhull}(\{P(\tau)\}_{\tau\in\mathcal{T}})$triangleright Returns the indices on the hull of all $P(\tau)$.
@@
13:         $maxSize\leftarrow\lfloor\alpha|\mathcal{T}|+\beta\rfloor$
14:     end if
15:end while
16:return $n$ and $logLikelihoodRatio$
Empirical Pruning and Run time Evaluations
@@
We give a link to an implementation of Algorithm 1 for a change in the mean of $p$-dimensional Gaussian signal in about a hundred lines of R code in the Supplementary Material. Our code is based on the convhulln function implementing the QuickHull algorithm in the geometry R package [59] and for $p=3$ it is able to process in 2 minutes $n=5\times 10^{5}$ data points on an Intel(R) Xeon(R) Gold 6252 CPU @ 2.10GHz processor choosing $\alpha=2$ and $\beta=1$. As can be seen in Figure 5 (left), such run time is comparable to that of the ocd R package [38].
@@
We also implemented Algorithm 1 for a change in the mean of multi-dimensional Gaussian signal (see Table 1) in R/C++ using the qhull library (see link in Supplementary Material). We evaluated its running time. We simulated data as in the empirical validation of Theorem 6 on the expected number of faces and vertices for dimension $p=1$ to $6$. The average run times as a function of $n$ are presented in the log-scale in Figure 4. As expected, it is faster than our R implementation: for example for $p=3$ it processes in 2 minutes $n=10^{6}$ rather than $5\times 10^{5}$. Note also that the run times for a known or unknown $\eta_{1}$ are almost identical. Finally, we made a simple linear regression to model the logarithm of the run time as a function of $\log(n)$. The estimated slope coefficients are reported in Figure 4 and are all smaller than $1.2$.
@@
Figure 4: Run times in seconds of MdFOCuS with known (dashed line) and unknown (full line) pre-change parameter $\eta_{1}$, both with $\alpha=2$ and $\beta=1$, in dimension $p=1,\dots,6$ using time series $x_{1:n}$ (without change, $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$ i.i.d). Run times are averaged over $100$ data sets. We considered a maximum running time of $20$ minutes (horizontal dotted black line) for the algorithms. That is why we do not report run times for $p=4,5$ and $6$ for large number of observations. We report here the slope of the curves estimated using a simple linear regression: for known $\eta_{1}$: $\approx 1.005$ ($p=1$), $\approx 1.116$ ($p=2$), $\approx 1.180$ ($p=3$), $\approx 1.196$ ($p=4$), $\approx 1.143$ ($p=5$) and $\approx 1.202$ ($p=6$); for unknown $\eta_{1}$: $\approx 1.006$ ($p=1$), $\approx 1.118$ ($p=2$), $\approx 1.183$ ($p=3$), $\approx 1.202$ ($p=4$), $\approx 1.144$ ($p=5$) and $\approx 1.197$ ($p=6$).
4.4 MdFOCuS(0) dyadic, MdFOCuS(0) Sp and MdFOCuS(0) Sp $\tilde{p}$d : Three Variants of the Algorithm
In this subsection, we describe three variants of the previous algorithm. The first, MdFOCuS(0) dyadic, is slower but provably quasi-linear. The second, MdFOCuS(0) Sp, computes sparse likelihood ratio statistics. The third, MdFOCuS(0) $\tilde{p}d$ Sp, is an approximation designed to handle large values of $p$.
4.4.1 MdFOCuS(0) dyadic: a more complex yet provably quasi-linear online algorithm
@@
Algorithm 1 is exact and empirically fast but it is non-trivial to derive its worst or expected time complexity as the time at which we run QuickHull is stochastic. In Supplementary Material F we propose Algorithm 3 that is closely related and easier to analyse theoretically. This algorithm updates the set $\mathcal{T}$ (lines 11 to 14 in Algorithm 1) in a deterministic manner on predetermined subsets or chunks of the observations. This deterministic choice of chunks is likely improvable but it allows for a precise quantification of the expected complexity of the algorithm under the assumption of Theorem 6 (or Corollary 1) and under the assumption that the convex hull algorithm is at worst quadratic. To be specific here are our assumptions.
@@
This is true under the conditions of Corollary 1.
@@
We assume that the convex hull algorithm we are using is at worst quadratic in time and assume there exist two positive constants $c_{3}$ and $c_{4}$ such that for $n$ points the time complexity is $Time(n)\leq c_{3}n^{2}+c_{4}$.
The assumption on the quadratic complexity is true for QuickHull in dimensions 2 and 3 (so for $p\leq 2$) [48, 63] and using the algorithm of Chazelle [64] it is also true in dimension 4 (so for $p\leq 3$). Still following Chazelle this is no longer true for any dimension strictly larger than $4$ and therefore $p>3$.
In essence, our deterministic Algorithm 3 computes the convex hull using a divide and conquer strategy. Out of simplicity let us consider the offline setting and assume $n-1$ is a power of $2$, that is $n-1=2^{\log_{2}(n-1)}.$ Rather than running the hull algorithm on all points $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$ we first run the algorithm on smaller chunks of size $2^{q_{min}}$ : $\tau\in\{2^{q_{min}}j+1,\ldots,2^{q_{min}}j+2^{q_{min}}\},$ with $q_{min}$ a constant chosen by the user and $j$ a number smaller than $(n-1)/2^{q_{min}}$, and then dyadically merge the results. To be specific for each scale $q$ with $q_{min}<q<\log_{2}(n-1)$ we get the point on chunk $\{2^{q+1}j+1,\ldots,2^{q+1}j+2^{q+1}\}$, by merging the results of the two smaller chunks $\{2^{q+1}j+1,\ldots,2^{q+1}j+2^{q}\},$ and $\{2^{q+1}j+2^{q}+1,\ldots,2^{q+1}j+2^{q+1}\}.$ This merging is computationally efficient because, under Assumption 1, the number of points on the hull of a chunk of size $2^{q}$ is expected to be small (of order $(\log(2^{q}))^{p}=(q\log(2))^{p}$).
One difficulty in Algorithm 3 is to make this dyadic or divide and conquer merging in an online fashion. To the best of our efforts, this makes the formal description a bit tedious, and we refer the interested reader to the Supplementary Material F for a precise description. We now state our bound on the expected complexity of this algorithm.
@@
Under the assumption of Corollary 1 (see Assumption 1) and Assumption 2, the expected time complexity of Algorithm 2 is

### v3

@@ 4.3 MdFOCuS: a Simple Online Algorithm
In this section, we propose an algorithm to compute the GLR statistics at every time step $n$ for a known and unknown pre-change parameter $\eta_{1}$. In both cases, based on Theorem 5 we only need to consider the changepoints $\tau$ such that $P(\tau)$ is on the hull of $\left\{P(\tau^{\prime})\right\}_{\tau^{\prime}\in\{1,\dots,n-1\}}.$
@@
Let us call $\mathcal{T}_{n}$ the set of index $\tau$ such that $P(\tau)=(\tau,\sum_{t=1}^{\tau}x_{t})$ is on the convex hull of all $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$. Restating Theorem 4, the changepoint $\hat{\tau}$ maximising the likelihood (5) is in $\mathcal{T}_{n}$. Hence, at step $n$, we only need to compute the likelihood (and likelihood ratio statistics) of changepoints $\tau$ in $\mathcal{T}_{n}$. Furthermore, as already explained after Theorem 5 we can prune all $\tau$ that are not in $\mathcal{T}_{n}$ as they will never be in $\mathcal{T}_{n^{\prime}}$ for any $n^{\prime}\geq n$.
@@
Computing the convex hull from scratch at every time step $n$ is inefficient: it requires at least $\mathcal{O}(n)$ operations and leads to an overall complexity at least quadratic in $n$. Ideally, we would like to update the hull using an efficient online or dynamic convex hull algorithm as the one in Clarkson and Shor (1989). There exist implementations of such an algorithm in dimensions $p=2$ Melkman (1987) and $p=3$ Hert and Schirra (2023). However, for simplicity, here we rely on an iterative use of the offline QuickHull algorithm Barber et al. (1996) that also works for larger dimensions and is available in the geometry R package Habel et al. (2023).
@@
In detail, for any $n^{\prime}>n$ we have $\mathcal{T}_{n^{\prime}}\subset\mathcal{T}_{n}\cup\{n,\ldots,n^{\prime}-1\}$. Therefore if we have already computed $\mathcal{T}_{n}$, we only apply QuickHull to the points $P(\tau)$ with $\tau$ in $\mathcal{T}_{n}\cup\{n,\ldots,n^{\prime}-1\}$. A first idea would be to apply the QuickHull algorithm at every time step $n,n+1,\ldots$, that is compute $\mathcal{T}_{n}$ as the hull of all $P(\tau)$ with $\tau$ in $\mathcal{T}_{n-1}\cup\{n\}$. In practice, we discovered that running the QuickHull algorithm intermittently, rather than at every iteration, resulted in faster run times. Specifically, we maintain a set of linearly increasing indices, denoted as $\mathcal{T}^{\prime}_{n}$, and only periodically execute QuickHull to reconstruct the true $\mathcal{T}_{n}$. More precisely, to decide when to run QuickHull, we update a maximum size variable for $\mathcal{T}^{\prime}_{n}$ (denoted maxSize). Then, at step $n+1$:
  • if $|\mathcal{T}^{\prime}_{n}|$, is smaller than maxSize we set $\mathcal{T}^{\prime}_{n+1}$ to $\mathcal{T}^{\prime}_{n}\cup\{n\}$,
  • otherwise we compute $\mathcal{T}^{\prime}_{n+1}$ as the convex hull of all $P(\tau)$ with $\tau$ in $\mathcal{T}^{\prime}_{n}\cup\{n\}$.
@@
In the latter case we also update maxSize to $\texttt{maxSize}=\lfloor\alpha|\mathcal{T}^{\prime}_{n+1}|+\beta\rfloor$, with $\alpha\geq 1$ and $\beta\geq 0$. It is important to note that, in this way, we always retain a few more possible changepoints than necessary: $\mathcal{T}_{n+1}\subseteq\mathcal{T}^{\prime}_{n+1}$.
@@
If $\beta=1$ and $\alpha=1$ we run the QuickHull algorithm at every time step. Larger values of $\alpha$ mean we run it less often, but at the cost of having a longer list of indices in $\mathcal{T}^{\prime}_{n}$. We empirically evaluated different values of $\alpha$ with $\beta=1$, and found $\alpha=2$ gave good practical performance; see Supplementary Material F.2.
@@
A formal description of our procedure, called MdFOCuS is presented in Algorithm 1. The algorithm covers both known and unknown pre-change parameters $\eta_{1}$, with the only difference being the calculation of the likelihood (lines 7 and 8). Therefore, the two variants store the same number of candidate changepoints and have the same time and memory complexities. In detail, computing the full likelihood ratio statistics for any particular changepoint has a time complexity of $\mathcal{O}(p)$, and therefore lines 7 and 8 in Algorithm 1 have a complexity of $\mathcal{O}(|\mathcal{T}|p)$. Recovering the maximum has complexity $\mathcal{O}(|\mathcal{T}|)$. Overall, this is $\mathcal{O}(|\mathcal{T}|p)$.
Algorithm 1 MdFOCuS algorithm
1:Inputs: $\{x_{t}\}_{t=1,2,\dots}$ $p$-variate independent time series; $thrs$ threshold; $\eta_{1}$ pre-change parameter (if known); $\texttt{maxSize}>p+1$, $\alpha\geq 1$, $\beta\geq 0$, pruning parameters.
2:Outputs: stopping time $n$ and maximum likelihood change : $\hat{\tau}(.)$ or $\hat{\tau}(\eta_{1})$
3:$GLR\leftarrow-\infty,\quad\mathcal{T}\leftarrow\emptyset,\quad n\leftarrow 1$
4:while ($GLR<thrs$do
5:     $n\leftarrow n+1$
6:     $\mathcal{T}\leftarrow\mathcal{T}\cup\{n-1\}$
7:     $\hat{m}\leftarrow\max_{\tau\in\mathcal{T}}\{\max_{\eta_{1},\eta_{2}}\ell_{\tau,n}(\eta_{1},\eta_{2})\}$triangleright $\eta_{1}$ is fixed if pre-change known
8:     $GLR\leftarrow\hat{m}-\max_{\eta_{1},\eta_{2}}\ell_{n,n}(\eta_{1},\eta_{2})$triangleright $\eta_{1}$ is fixed if pre-change known
[…]
@@
10:         $\mathcal{T}\leftarrow\textsc{QuickHull}(\{P(\tau)\}_{\tau\in\mathcal{T}})$triangleright Find the indices on the hull of all $P(\tau)$
@@
11:         $\texttt{maxSize}\leftarrow\lfloor\alpha|\mathcal{T}|+\beta\rfloor$
12:     end if
13:end while
14:return $n$ and $GLR$
@@
A link to a simple R implementation of Algorithm 1 for a change in the mean of $p$-dimensional Gaussian signal can be found in the Supplementary Materials. Our code is based on the convhulln function implementing the QuickHull algorithm in the geometry R package, and for $p=3$ it takes 2 minutes to process $n=5\times 10^{5}$ data points on an Intel(R) Xeon(R) Gold 6252 CPU @ 2.10GHz processor choosing $\alpha=2$ and $\beta=1$. As can be seen in Figure 4 (left), such run time is comparable to that of the ocd R package Chen et al. (2022).
@@
We also implemented Algorithm 1 for a change in the mean of a multi-dimensional Gaussian signal in R/C++ using the qhull library. We use the Gaussian distribution as an example of a distribution with continuous density and independent dimensions. To be specific, for $p$ from $1$ to $6$, we simulated time series $x_{1:n}$, with $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$ i.i.d. We varied the time series length from $2^{10}+1(\approx 10^{3})$ to $2^{23}+1(\approx 8\times 10^{6})$ and simulated $100$ data sets for each length. The average run times as a function of $n$ are presented on the log-scale in Figure 3. As expected, it is faster than our R implementation: for example, for $p=3$ it processes in 2 minutes for $n=10^{6}$ rather than $5\times 10^{5}$. The run times for a known or unknown $\eta_{1}$ are almost identical. Finally, we made a simple linear regression to model the logarithm of the run time as a function of $\log(n)$. The estimated slope coefficients are reported in Figure 3 and are all smaller than $1.2$.
@@
Figure 3: Run times in seconds of MdFOCuS with known (dashed line) and unknown (full line) pre-change parameter $\eta_{1}$, both with $\alpha=2$ and $\beta=1$, in dimension $p=1,\dots,6$ for i.i.d Gaussian data. Run times are averaged over $100$ data sets. We considered a maximum running time of $20$ minutes (horizontal dotted black line) for the algorithms, and thus do not report run times for $p\geq 4$ for large $n$. The slope of the curves estimated are for known $\eta_{1}$: $\approx 1.005$ ($p=1$), $\approx 1.116$ ($p=2$), $\approx 1.180$ ($p=3$), $\approx 1.196$ ($p=4$), $\approx 1.143$ ($p=5$) and $\approx 1.202$ ($p=6$); and for unknown $\eta_{1}$: $\approx 1.006$ ($p=1$), $\approx 1.118$ ($p=2$), $\approx 1.183$ ($p=3$), $\approx 1.202$ ($p=4$), $\approx 1.144$ ($p=5$) and $\approx 1.197$ ($p=6$).
4.4 A provably quasi-linear online implementation
@@
We showed how Algorithm 1 is exact and empirically fast, but it is non-trivial to derive its worst or expected time complexity as the time at which we run QuickHull is stochastic. In Supplementary Material G we propose Algorithm 3 that is closely related but easier to analyse. This algorithm updates the set $\mathcal{T}$ in a deterministic manner on predetermined chunks of the observations. This deterministic choice of chunks allows for a precise quantification of the expected complexity of the algorithm using Corollary 1 and the assumption that the convex hull algorithm is at worst quadratic. To be specific, here are our assumptions.
@@
This is true under the conditions of Theorem 6 and Corollary 1. Next, define $Time(n)$ to be the worst-case computational cost of finding the convex hull of $n$ points.
@@
For our convex hull algorithm there exist two positive constants $c_{3}$ and $c_{4}$ such that $Time(n)\leq c_{3}n^{2}+c_{4}$.
The assumption on the quadratic complexity is true for QuickHull for $p\leq 2$ Barber et al. (1996) and using the algorithm of Chazelle (1993) it is true for $p\leq 3$.
In essence, Algorithm 3 computes the convex hull using a divide-and-conquer strategy. Out of simplicity, let us consider the offline setting and assume $n-1$ is a power of $2$, that is $n-1=2^{\log_{2}(n-1)}$. Rather than running the hull algorithm on all points $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$ we first run the algorithm on smaller chunks of size $2^{q_{min}}$ : $\tau\in\{2^{q_{min}}j+1,\ldots,2^{q_{min}}j+2^{q_{min}}\},$ with $q_{min}$ a constant chosen by the user and $j$ a number smaller than $(n-1)/2^{q_{min}}$, and then dyadically merge the results. To be specific for each scale $q$ with $q_{min}<q<\log_{2}(n-1)$ we get the points on chunk $\{2^{q+1}j+1,\ldots,2^{q+1}j+2^{q+1}\}$, by merging the results of the two smaller chunks $\{2^{q+1}j+1,\ldots,2^{q+1}j+2^{q}\},$ and $\{2^{q+1}j+2^{q}+1,\ldots,2^{q+1}j+2^{q+1}\}$. This merging is computationally efficient because, under Assumption 1, the number of points on the hull of a chunk of size $2^{q}$ is expected to be small (of order $(\log(2^{q}))^{p}=(q\log(2))^{p}$).
One difficulty in Algorithm 3 is to make this dyadic or divide-and-conquer merging in an online fashion. To the best of our efforts, this makes the formal description a bit tedious, and we refer the interested reader to the Supplementary Material G and Algorithm 2 and 3 for a precise description. We now state our bound on the expected complexity of Algorithms 2 and 3.
@@
Under Assumptions 1 and 2, the expected time complexity of Algorithm 2 and 3 is $\mathcal{O}\left(n\log(n)^{p+1}\right)$.
We empirically study the complexity of Algorithm 3 in Figure 16 of Supplementary Material G, where we see that it is slower than Algorithm 1. It is interesting to compare the bound on the computational cost with that of an oracle algorithm that knows the set of points on the convex hull at each iteration and whose computational cost is only evaluating the test statistic for each point on the hull. This would have a complexity of order $n\log^{p}(n)$, as $\sum_{t=1}^{n}\log^{p}(t)\geq n/2\log^{p}(n/2)$. This is just a factor of $\log n$ faster than the bound in Lemma 1.
4.5 Two variant of MdFOCuS
In this subsection, we describe two variants of the algorithm, arising from variations in (i) the test statistic construction and (ii) the convex hull update. These variants are available regardless of whether or not the pre-change parameter is known.
The first variant will allow MdFOCuS to improve on scenarios where only a subset of the coordinates change: out of simplicity in lines 7 and 8 of Algorithm 1 we only consider the likelihood ratio statistics (5) that assumes all coordinates are changing. However, based on Theorem 5, the changepoints maximising the sparse likelihood ratio statistics (2), where only $s$ coordinates are changing, also correspond to points on the hull, and we can compute the test statistics in this case by replacing lines 7 and 8 with the corresponding sparse likelihoods.
To compute all sparse max likelihood ratio statistics we sort the per-dimension ratio statistics (this can be done in $\mathcal{O}(p\log(p))$) and sum these from the largest to the smallest in $\mathcal{O}(p)$ to iteratively recover the best for all sparsity levels $s$. We then recover the maximum for each sparsity level in $\mathcal{O}(\mathcal{T})$. If we consider all sparsity levels the total complexity is $\mathcal{O}(|\mathcal{T}|p\log(p))$.
In practice, it makes sense to consider a subset of all sparsity levels. In our simulations we consider: $s\in\{2^{0},2^{1},2^{2},\ldots,2^{\lfloor\log_{2}(p)\rfloor},p\}$. In a similar fashion to the FOCuS approximation, we also consider the sum of the max likelihood ratio statistics for each coordinate. As can be seen in Figure 4 , the empirical run time of MdFOCuS with sparse statistics is comparable to that of ocd Chen et al. (2022).
The second variant extends the algorithm to high-dimensional settings, where a full convex hull update is computationally prohibitive. In fact, under the condition of Theorem 6, Algorithm 1 is expected to store $\mathcal{O}(\log^{p}(n))$ candidates and the $\log^{p}(n)$ term means that the algorithm is impractical for $p$ larger than $5,$ as exemplified in Figure 3. In Supplementary Material H, we propose an approximate algorithm by recovering only a subset of the points on the hull. We obtain this subset by considering projections on at most $\tilde{p}$ variates, running the hull algorithm on each projection, and taking the union of the time-points (of the vertices) retained for each projection. For each projection or group of variates, under the condition of Theorem 6, we store $\mathcal{O}(\log^{\tilde{p}}(n))$ candidates. Considering $O(p)$ such groups we store $\mathcal{O}(p\log^{\tilde{p}}(n))$ candidates, thus reducing the computational complexity.
This heuristic for $\tilde{p}=2$, implemented in R, runs on average in about 100 seconds for $p=100$ and $n=10000$. We further explore the run times of this heuristic in Figure 4 and show that its run time is comparable to ocd. As shown in Section 4.7 it provides good statistical performance.